OpenStudy (anonymous):

Identify as a direct variation, inverse variation or neither. y = 6x

OpenStudy (anonymous):

WILL METAL

Nnesha (nnesha):

what is metal ? i want chocolates ...

Nnesha (nnesha):

equation for direct variation is y = kx ( k is multiplying with x) inverse variation y = k/x <-- division( k is divided by x)

Nnesha (nnesha):

so y = 6x is it direct variation or inverse ?

OpenStudy (anonymous):

x and Y are directly proportional to each other since their division remains constant. y=6x y/x=6 As value of one of the variables increases, the value of the other is bound to increase to keep the quotient constant to 6.

OpenStudy (anonymous):

your chocolate cx lol and is it inverse?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

inverse ?? how ?

Nnesha (nnesha):

that chocolate >.< :/

OpenStudy (anonymous):

then it is direct cx lol and what don't it look yummy xD

Nnesha (nnesha):

inverse ? how ? direct ? how ? what do you think it's direct ?

OpenStudy (anonymous):

im actully lost xc thats why lol

Nnesha (nnesha):

alright y = 6x in this equation k = 6 which is constant of variation if k is multiplying with x then it's direct if k is dividing by x then inverse inverse ( division k/x) direct(multiplication kx)

OpenStudy (anonymous):

so it is direct thanks cx ^.^

Nnesha (nnesha):

what do you thin its direct ?? explain ? o^_^o